# Selmer stability for elliptic curves in Galois $\ell$-extensions

### v1

@@ 2.3. The method of Reichardt and Scholz
Let $G$ be an $\ell$-group with #$\#G=\ell^{n}$. Let $L/\mathbb{Q}$ be a Galois extension with $\operatorname{Gal}(L/\mathbb{Q})\simeq G$ and choose $N\geq n$. The extension $L/\mathbb{Q}$ satisfies the Scholz property for $N$, denoted $(\mathfrak{S}_{N})$, if:
@@
Let $\ell$ be an odd prime number, $G$ be a finite group with #$\#G=\ell^{n}$, and fix $N\geq n$. Suppose that there exists a Galois extension $L/\mathbb{Q}$ with $\operatorname{Gal}(L/\mathbb{Q})\simeq G$ which satisfies $(\mathfrak{S}_{N})$. Let $\{p_{1},\dots,p_{m}\}$ denote the primes that ramify in $L$. Then the embedding problem for $L$ and $\widetilde{G}$ is solvable. Moreover, the solution $\widetilde{L}$ can be chosen to satisfy $(\mathfrak{S}_{N})$ with at most one additional prime outside $\{p_{1},\dots,p_{m}\}$ being ramified in $\widetilde{L}$.
@@
[…]
@@
$\delta(\mathcal{P}_{L}):=\frac{1}{\left[L\left(\mu_{\ell^{N}},\sqrt[\ell]{p_{1}},\ldots,\sqrt[\ell]{p_{m}}\right):\mathbb{Q}\right]}.$
Let $q\in\mathcal{P}_{L}$ and $M_{q}\subset\mathbb{Q}(\mu_{q})$ be the field such that $\operatorname{Gal}(M_{q}/\mathbb{Q})\simeq\mathbb{Z}/\ell\mathbb{Z}$. Set $\widetilde{L}:=L\cdot M_{q}$. Since $q$ is totally ramified in $M_{q}$ and split in $L$, it follows that $L\cap M_{q}=\mathbb{Q}$, therefore, $\operatorname{Gal}(\widetilde{L}/\mathbb{Q})\simeq G\times\mathbb{Z}/\ell\mathbb{Z}=\widetilde{G}$. It is easy to see that $\widetilde{L}$ satisfies $(\mathfrak{S}_{N})$. In greater detail, each of the primes $p_{i}$ are of the form $p_{i}\equiv 1\bmod\ell^{N}$ as $L$ satisfies ($\mathfrak{S}_{N}$). The prime $q$ is assumed to satisfy $q\equiv 1\bmod\ell^{N}$. Thus, all primes that ramify in $\widetilde{L}$ are $\equiv 1\bmod\ell^{N}$. Given a prime $p_{i}$ for $i=1,\cdots,m$, let $\widetilde{v}$ be a prime of $\widetilde{L}$ that lies above $p_{i}$. Let $v$ (resp. $v^{\prime}$) be a prime of $L$ (resp. $M_{q}$) such that $\widetilde{v}|v$ (resp. $\widetilde{v}|v^{\prime}$), as depicted below:
@@
Since $L$ satisfies $(\mathfrak{S}_{N})$, we have that $f(v\mid p_{i})=1$. By condition (iii) in the choice of $\mathcal{P}_{L}$, $p_{i}$ splits completely in $M_{q}$. Therefore, we deduce that $f(\widetilde{v}\mid p_{i})=1$, i.e., $\widetilde{L}_{\widetilde{v}}/\mathbb{Q}_{p_{i}}$ is totally ramified. Since $q$ splits completely in $L$ and is totally ramified in $M_{q}$, $f(\widetilde{w}\mid q)=1$ for any prime $\widetilde{w}$ of $\widetilde{L}$ that lies above $q$. Thus, $\widetilde{L}$ as defined above is a solution the embedding problem and satisfies $(\mathfrak{S}_{N})$. The set of primes ramified in $\widetilde{L}$ is precisely $\{p_{1},\cdots,p_{m}$, $q\}$ and therefore, $\widetilde{L}$ is ramified at exactly one additional prime.
@@
Recall that $L$ induces a surjective homomorphism $\varphi:\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow G$ and that there is a bijection between the central extensions $\widetilde{G}$ of $G$ by $\mathbb{Z}/\ell\mathbb{Z}$ and the elements of $H^{2}\left(G,\mathbb{Z}/\ell\mathbb{Z}\right)$. Let $\theta=\theta_{\widetilde{G}}\in H^{2}\left(G,\mathbb{Z}/\ell\mathbb{Z}\right)$ be the element corresponding to the extension $\widetilde{G}$. The map $\varphi$ induces a homomorphism
$\varphi^{*}:H^{2}\left(G,\mathbb{Z}/\ell\mathbb{Z}\right)\longrightarrow H^{2}\left(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z}\right).$ (2.3)
@@
Let $\mathcal{G}$, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be groups, and let $f_{1}:\mathcal{G}_{1}\to\mathcal{G}$ and $f_{2}:\mathcal{G}_{2}\to\mathcal{G}$ be surjective homomorphisms. The fibre product $\mathcal{G}_{1}\times_{\mathcal{G}}\mathcal{G}_{2}$ is the group
@@
with projections $\pi_{1}:\mathcal{G}_{1}\times_{\mathcal{G}}\mathcal{G}_{2}\to\mathcal{G}_{1}$ and $\pi_{2}:\mathcal{G}_{1}\times_{\mathcal{G}}\mathcal{G}_{2}\to\mathcal{G}_{2}$. It satisfies the universal property: for any group $H$ with homomorphisms $q_{1}:H\to\mathcal{G}_{1}$ and $q_{2}:H\to\mathcal{G}_{2}$ such that $f_{1}\circ q_{1}=f_{2}\circ q_{2}$, there exists a unique homomorphism $q:H\to\mathcal{G}_{1}\times_{\mathcal{G}}\mathcal{G}_{2}$ making the diagram below commute:
${H}$${\mathcal{G}_{1}\times_{\mathcal{G}}\mathcal{G}_{2}}$${\mathcal{G}_{1}}$${\mathcal{G}_{2}}$${\mathcal{G}.}$$\scriptstyle{q_{1}}$$\scriptstyle{q_{2}}$$\scriptstyle{q}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{f_{1}}$$\scriptstyle{f_{2}}$
@@
This is a well known result, we give a sketch of the argument. Setting $\mathfrak{G}:=\widetilde{G}\times_{G}\operatorname{G}_{\mathbb{Q}},$ consider the fiber product diagram:
@@
It is easy to see that the top row is a central extension corresponding to $\varphi^{*}(\theta)$. This sequence splits if and only if $\varphi^{*}(\theta)=0$. If the extension splits, there exists a section $j:\operatorname{G}_{\mathbb{Q}}\hookrightarrow\mathfrak{G}$ to the projection map $\pi_{2}:\mathfrak{G}\twoheadrightarrow\operatorname{G}_{\mathbb{Q}}$. Then $\widetilde{\varphi}:=p_{1}\circ j$ is a lift of $\varphi$. Since it is assumed that (2.2) is nonsplit, $\widetilde{\varphi}$ is surjective. The proof of the converse follows along similar lines, and is omitted. ∎
In what follows, for ease of notation, set $H^{i}(F,\cdot):=H^{i}(\operatorname{G}_{F},\cdot)$, where $F$ is a field of characteristic 00. Given a prime number $p$, denote by
$\operatorname{res}^{i}_{p}:H^{i}(\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})\longrightarrow H^{i}(\mathbb{Q}_{p},\mathbb{Z}/\ell\mathbb{Z})$
@@
$\alpha:H^{2}(\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})\xrightarrow{\bigoplus_{p}\operatorname{res}_{p}^{2}}\bigoplus_{p}H^{2}(\mathbb{Q}_{p},\mathbb{Z}/\ell\mathbb{Z}).$
@@
$\alpha_{K}:H^{2}(K,\mathbb{Z}/\ell\mathbb{Z})\longrightarrow\bigoplus_{v\in M_{K}}H^{2}(K_{v},\mathbb{Z}/\ell\mathbb{Z})$
@@
Given a prime $p$, let $G_{p}:=\operatorname{Gal}(L_{v}/\mathbb{Q}_{p})\subset G$, where $v$ is a prime of $L$ lying over $p$. A different choice of $v$ gives rise to the same subgroup of $G$ up to conjugation by an element of $G$. Set $\theta_{p}$ to denote the image of $\theta$ with respect to the natural restriction map:
$\operatorname{res}_{p}:H^{2}(G,\mathbb{Z}/\ell\mathbb{Z})\to H^{2}(G_{p},\mathbb{Z}/\ell\mathbb{Z}).$
@@
and let $\varphi_{p}:\operatorname{G}_{\mathbb{Q}_{p}}\twoheadrightarrow G_{p}$ be the natural quotient map. The local embedding problem then asks whether $\varphi_{p}$ admits a lift to a homomorphism $\widetilde{\varphi}_{p}:\operatorname{G}_{\mathbb{Q}_{p}}\rightarrow\widetilde{G}_{p}$. Here we do not insist that $\widetilde{\varphi}_{p}$ is surjective. By arguments similar to those in the proof of Proposition 2.5, we find that the local embedding problem at $p$ has a solution if and only if $\varphi_{p}^{*}(\theta_{p})=0$.
@@
With the notation as above, a surjective lift $\widetilde{\varphi}:\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow\widetilde{G}$ exists if and only if the local embedding problem is solvable for all primes $p$.
@@
By Proposition 2.5, a surjective lift $\widetilde{\varphi}$ exists if and only if $\varphi^{*}(\theta)=0$. The map $\alpha:H^{2}(\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})\rightarrow\bigoplus_{p\in M_{\mathbb{Q}}}H^{2}(\mathbb{Q}_{p},\mathbb{Z}/\ell\mathbb{Z})$ maps $\varphi^{*}(\theta)$ to the tuple $\alpha(\varphi^{*}(\theta))=\left(\varphi_{p}^{*}(\theta_{p})\right)_{p}$. Proposition 2.6 shows that the map $\alpha$ is injective, so it suffices to verify that $\varphi_{p}^{*}(\theta_{p})=0$ for all primes $p$. This is equivalent to the existence of a local lift $\widetilde{\varphi}_{p}:\operatorname{G}_{\mathbb{Q}_{p}}\to\widetilde{G}_{p}$ at each prime $p$, which establishes the claim. ∎
We now show that the local embedding problem is solvable. We then deduce from Proposition 2.7 that a surjective lift $\widetilde{\varphi}:\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow\widetilde{G}$ of $\varphi:\operatorname{G}_{\mathbb{Q}}\rightarrow G$ does exist. Given a finite group $\mathcal{G}$, recall that the Frattini subgroup $\Phi(\mathcal{G})$ is the intersection of all proper maximal subgroups of $\mathcal{G}$. If $\mathcal{G}$ is an $\ell$-group, then $\Phi(\mathcal{G})=\mathcal{G}^{p}[\mathcal{G},\mathcal{G}]$, and $\mathcal{G}/\Phi(\mathcal{G})\simeq(\mathbb{Z}/\ell\mathbb{Z})^{d}$ for some $d\geq 0$. It is well known that if $\mathcal{G}$ is a finite $\ell$-group, then $\mathcal{G}$ is cyclic if and only if the Frattini quotient $\mathcal{G}/\Phi(\mathcal{G})\simeq\mathbb{Z}/\ell\mathbb{Z}$.
@@
Suppose $L/\mathbb{Q}$ is a Galois extension with $\operatorname{Gal}(L/\mathbb{Q})\simeq G$ and canonical surjection $\varphi:\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow G$ as before. Then, for every prime $p$, the local embedding problem (2.4) is solvable.
@@
First suppose that $p$ is unramified in $L$. In this case, $G_{p}$ is cyclic, generated by the Frobenius element. Let $\mathbb{Q}_{p}^{\operatorname{nr}}$ be the maximal unramified extension of $\mathbb{Q}_{p}$ and note that $\operatorname{Gal}(\mathbb{Q}_{p}^{\operatorname{nr}}/\mathbb{Q}_{p})$ is isomorphic to $\widehat{\mathbb{Z}}$. We seek a lift $\widetilde{\varphi}_{p}:\widehat{\mathbb{Z}}\rightarrow\widetilde{G}_{p}$ of $\varphi_{p}$. In order to show that $\widetilde{\varphi}_{p}$ exists, it suffices to prove that $\widetilde{G}_{p}$ is cyclic. Indeed, $G_{p}$ is cyclic, in particular, abelian. Hence $\mathbb{Z}/\ell\mathbb{Z}$ contains $[\widetilde{G}_{p},\widetilde{G}_{p}]$. If $\mathbb{Z}/\ell\mathbb{Z}=[\widetilde{G}_{p},\widetilde{G}_{p}]$, then, $\mathbb{Z}/\ell\mathbb{Z}$ is contained in $\Phi(\widetilde{G}_{p})=[\widetilde{G}_{p},\widetilde{G}_{p}]\widetilde{G}_{p}^{\ell}$. Therefore, there is a surjection $\widetilde{G}_{p}/\left(\mathbb{Z}/\ell\mathbb{Z}\right)\twoheadrightarrow\widetilde{G}_{p}/\Phi(\widetilde{G}_{p})$. In particular, we find that $\widetilde{G}_{p}/\Phi(\widetilde{G}_{p})$ must be cyclic. It follows that $\widetilde{G}_{p}$ must be cyclic as well. This contradicts the assumption that $[\widetilde{G}_{p},\widetilde{G}_{p}]=\mathbb{Z}/\ell\mathbb{Z}$. Thus, $\widetilde{G}_{p}$ is abelian. Since $G_{p}$ is cyclic and the sequence (2.4) is non-split, this forces $\widetilde{G}_{p}$ to be cyclic. Therefore, $\varphi_{p}:\widehat{\mathbb{Z}}\rightarrow G_{p}$ lifts to a map $\widetilde{\varphi}_{p}:\widehat{\mathbb{Z}}\rightarrow\widetilde{G}_{p}$.

### v2

@@ 2.3. The method of Reichardt and Scholz
Let $G$ be an $\ell$-group with $G=\ell^{n}$. Let $L/\mathbb{Q}$ be a Galois extension with $\operatorname{Gal}(L/\mathbb{Q})\simeq G$ and choose $N\geq n$. The extension $L/\mathbb{Q}$ satisfies the Scholz property for $N$, denoted $(\mathfrak{S}_{N})$, if:
@@
Let $\ell$ be an odd prime number, $G$ be a finite group with $G=\ell^{n}$, and fix $N\geq n$. Suppose that there exists a Galois extension $L/\mathbb{Q}$ with $\operatorname{Gal}(L/\mathbb{Q})\simeq G$ which satisfies $(\mathfrak{S}_{N})$. Let $\{p_{1},\dots,p_{m}\}$ denote the primes that ramify in $L$. Then the embedding problem for $L$ and $\widetilde{G}$ is solvable. Moreover, the solution $\widetilde{L}$ can be chosen to satisfy $(\mathfrak{S}_{N})$ with at most one additional prime outside $\{p_{1},\dots,p_{m}\}$ being ramified in $\widetilde{L}$.
@@
First, we consider the case when (2.2) is split. Let $p_{1},\dots,p_{m}$ be the rational primes which are ramified in $L$ and $q$ be a prime number. It is easy to see that $q$ splits completely in the field $L\left(\mu_{\ell^{N}},\sqrt[\ell]{p_{1}},\ldots,\sqrt[\ell]{p_{m}}\right)$ if and only if the following conditions are satisfied:
@@
$\delta(\mathcal{P}_{L})\mathrel{\mathop{\ordinarycolon}}=\frac{1}{\left[L\left(\mu_{\ell^{N}},\sqrt[\ell]{p_{1}},\ldots,\sqrt[\ell]{p_{m}}\right)\mathrel{\mathop{\ordinarycolon}}\mathbb{Q}\right]}$.
Let $q\in\mathcal{P}_{L}$ and $M_{q}\subset\mathbb{Q}(\mu_{q})$ be the field such that $\operatorname{Gal}(M_{q}/\mathbb{Q})\simeq\mathbb{Z}/\ell\mathbb{Z}$. Set $\widetilde{L}\mathrel{\mathop{\ordinarycolon}}=L\cdot M_{q}$. Since $q$ is totally ramified in $M_{q}$ and split in $L$, it follows that $L\cap M_{q}=\mathbb{Q}$, therefore, $\operatorname{Gal}(\widetilde{L}/\mathbb{Q})\simeq G\times\mathbb{Z}/\ell\mathbb{Z}=\widetilde{G}$. It is easy to see that $\widetilde{L}$ satisfies $(\mathfrak{S}_{N})$. In greater detail, each of the primes $p_{i}$ are of the form $p_{i}\equiv 1\bmod\ell^{N}$ as $L$ satisfies ($\mathfrak{S}_{N}$). The prime $q$ is assumed to satisfy $q\equiv 1\bmod\ell^{N}$. Thus, all primes that ramify in $\widetilde{L}$ are $\equiv 1\bmod\ell^{N}$. Given a prime $p_{i}$ for $i=1,\cdots,m$, let $\widetilde{v}$ be a prime of $\widetilde{L}$ that lies above $p_{i}$. Let $v$ (resp. $v^{\prime}$) be a prime of $L$ (resp. $M_{q}$) such that $\widetilde{v}|v$ (resp. $\widetilde{v}|v^{\prime}$), as depicted below:
@@
Since $L$ satisfies $(\mathfrak{S}_{N})$, we have that $f(v\mid p_{i})=1$. By condition (iii) in the choice of $\mathcal{P}_{L}$, $p_{i}$ splits completely in $M_{q}$. Therefore, we deduce that $f(\widetilde{v}\mid p_{i})=1$, i.e., $\widetilde{L}_{\widetilde{v}}/\mathbb{Q}_{p_{i}}$ is totally ramified. Since $q$ splits completely in $L$ and is totally ramified in $M_{q}$, $f(\widetilde{w}\mid q)=1$ for any prime $\widetilde{w}$ of $\widetilde{L}$ that lies above $q$. Thus, $\widetilde{L}$ as defined above is a solution the embedding problem and satisfies $(\mathfrak{S}_{N})$. The set of primes ramified in $\widetilde{L}$ is precisely $\{p_{1},\cdots,p_{m}$, q}q\} and therefore, $\widetilde{L}$ is ramified at exactly one additional prime.
@@
We note that step 1 corresponds to Proposition 2.9, step 2 corresponds to Proposition 2.13, and step 3 corresponds to Proposition 2.15. Recall that $L$ induces a surjective homomorphism $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow G$ and that there is a bijection between the central extensions $\widetilde{G}$ of $G$ by $\mathbb{Z}/\ell\mathbb{Z}$ and the elements of $H^{2}\left(G,\mathbb{Z}/\ell\mathbb{Z}\right)$. Let $\theta=\theta_{\widetilde{G}}\in H^{2}\left(G,\mathbb{Z}/\ell\mathbb{Z}\right)$ be the element corresponding to the extension $\widetilde{G}$. The map $\varphi$ induces a homomorphism
$\varphi^{*}\mathrel{\mathop{\ordinarycolon}}H^{2}\left(G,\mathbb{Z}/\ell\mathbb{Z}\right)\longrightarrow H^{2}\left(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z}\right).$ (2.3)
@@
Let $\mathcal{G}$, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be groups, and let $f_{1}\mathrel{\mathop{\ordinarycolon}}\mathcal{G}_{1}\to\mathcal{G}$ and $f_{2}\mathrel{\mathop{\ordinarycolon}}\mathcal{G}_{2}\to\mathcal{G}$ be surjective homomorphisms. The fibre product $\mathcal{G}_{1}\times_{\mathcal{G}}\mathcal{G}_{2}$ is the group
@@
with projections $\pi_{1}\mathrel{\mathop{\ordinarycolon}}\mathcal{G}_{1}\times_{\mathcal{G}}\mathcal{G}_{2}\to\mathcal{G}_{1}$ and $\pi_{2}\mathrel{\mathop{\ordinarycolon}}\mathcal{G}_{1}\times_{\mathcal{G}}\mathcal{G}_{2}\to\mathcal{G}_{2}$. It satisfies the universal property: for any group $H$ with homomorphisms $q_{1}\mathrel{\mathop{\ordinarycolon}}H\to\mathcal{G}_{1}$ and $q_{2}\mathrel{\mathop{\ordinarycolon}}H\to\mathcal{G}_{2}$ such that $f_{1}\circ q_{1}=f_{2}\circ q_{2}$, there exists a unique homomorphism $q\mathrel{\mathop{\ordinarycolon}}H\to\mathcal{G}_{1}\times_{\mathcal{G}}\mathcal{G}_{2}$ making the diagram below commute:
[…]
@@
This is a well known result, we give a sketch of the argument. Setting $\mathfrak{G}\mathrel{\mathop{\ordinarycolon}}=\widetilde{G}\times_{G}\operatorname{G}_{\mathbb{Q}},$ consider the fiber product diagram:
@@
Denote by $\pi_{1}\mathrel{\mathop{\ordinarycolon}}\mathfrak{G}\rightarrow\widetilde{G}$ and $\pi_{2}\mathrel{\mathop{\ordinarycolon}}\mathfrak{G}\rightarrow\operatorname{G}_{\mathbb{Q}}$ the projection maps to the first and second factors. It is easy to see that the top row is a central extension corresponding to $\varphi^{*}(\theta)$. This sequence splits if and only if $\varphi^{*}(\theta)=0$. If the extension splits, there exists a section $j\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\hookrightarrow\mathfrak{G}$ to the projection map $\pi_{2}\mathrel{\mathop{\ordinarycolon}}\mathfrak{G}\twoheadrightarrow\operatorname{G}_{\mathbb{Q}}$. Then $\widetilde{\varphi}\mathrel{\mathop{\ordinarycolon}}=\pi_{1}\circ j$ is a lift of $\varphi$. Since it is assumed that (2.2) is nonsplit, $\widetilde{\varphi}$ is surjective. The proof of the converse follows along similar lines, and is omitted. ∎
In what follows, for ease of notation, set $H^{i}(F,\cdot)\mathrel{\mathop{\ordinarycolon}}=H^{i}(\operatorname{G}_{F},\cdot)$, where $F$ is a field of characteristic 0. Given a prime number $p$, denote by
$\operatorname{res}^{i}_{p}\mathrel{\mathop{\ordinarycolon}}H^{i}(\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})\longrightarrow H^{i}(\mathbb{Q}_{p},\mathbb{Z}/\ell\mathbb{Z})$
@@
$\alpha\mathrel{\mathop{\ordinarycolon}}H^{2}(\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})\xrightarrow{\bigoplus_{p}\operatorname{res}_{p}^{2}}\bigoplus_{p}H^{2}(\mathbb{Q}_{p},\mathbb{Z}/\ell\mathbb{Z}).$
@@
$\alpha_{K}\mathrel{\mathop{\ordinarycolon}}H^{2}(K,\mathbb{Z}/\ell\mathbb{Z})\longrightarrow\bigoplus_{v\in M_{K}}H^{2}(K_{v},\mathbb{Z}/\ell\mathbb{Z})$
@@
Given a prime $p$, let $G_{p}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(L_{v}/\mathbb{Q}_{p})\subset G$, where $v$ is a prime of $L$ lying over $p$. A different choice of $v$ gives rise to the same subgroup of $G$ up to conjugation by an element of $G$. Set $\theta_{p}$ to denote the image of $\theta$ with respect to the natural restriction map:
$\operatorname{res}_{p}\mathrel{\mathop{\ordinarycolon}}H^{2}(G,\mathbb{Z}/\ell\mathbb{Z})\to H^{2}(G_{p},\mathbb{Z}/\ell\mathbb{Z}).$
@@
and let $\varphi_{p}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}_{p}}\twoheadrightarrow G_{p}$ be the natural quotient map. The local embedding problem then asks whether $\varphi_{p}$ admits a lift to a homomorphism $\widetilde{\varphi}_{p}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}_{p}}\rightarrow\widetilde{G}_{p}$. Here we do not insist that $\widetilde{\varphi}_{p}$ is surjective. By arguments similar to those in the proof of Proposition 2.5, we find that the local embedding problem at $p$ has a solution if and only if $\varphi_{p}^{*}(\theta_{p})=0$.
@@
With the notation as above, a surjective lift $\widetilde{\varphi}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow\widetilde{G}$ exists if and only if the local embedding problem is solvable for all primes $p$.
@@
By Proposition 2.5, a surjective lift $\widetilde{\varphi}$ exists if and only if $\varphi^{*}(\theta)=0$. The map $\alpha\mathrel{\mathop{\ordinarycolon}}H^{2}(\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})\rightarrow\bigoplus_{p\in M_{\mathbb{Q}}}H^{2}(\mathbb{Q}_{p},\mathbb{Z}/\ell\mathbb{Z})$ maps $\varphi^{*}(\theta)$ to the tuple $\alpha(\varphi^{*}(\theta))=\left(\varphi_{p}^{*}(\theta_{p})\right)_{p}$. Proposition 2.6 shows that the map $\alpha$ is injective, so it suffices to verify that $\varphi_{p}^{*}(\theta_{p})=0$ for all primes $p$. This is equivalent to the existence of a local lift $\widetilde{\varphi}_{p}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}_{p}}\to\widetilde{G}_{p}$ at each prime $p$, which establishes the claim. ∎
We now show that the local embedding problem is solvable. We then deduce from Proposition 2.7 that a surjective lift $\widetilde{\varphi}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow\widetilde{G}$ of $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow G$ does exist. Given a finite group $\mathcal{G}$, recall that the Frattini subgroup $\Phi(\mathcal{G})$ is the intersection of all proper maximal subgroups of $\mathcal{G}$. If $\mathcal{G}$ is an $\ell$-group, then $\Phi(\mathcal{G})=\mathcal{G}^{p}[\mathcal{G},\mathcal{G}]$, and $\mathcal{G}/\Phi(\mathcal{G})\simeq(\mathbb{Z}/\ell\mathbb{Z})^{d}$ for some $d\geq 0$. It is well known that if $\mathcal{G}$ is a finite $\ell$-group, then $\mathcal{G}$ is cyclic if and only if the Frattini quotient $\mathcal{G}/\Phi(\mathcal{G})\simeq\mathbb{Z}/\ell\mathbb{Z}$.
@@
Suppose $L/\mathbb{Q}$ is a Galois extension with $\operatorname{Gal}(L/\mathbb{Q})\simeq G$ and canonical surjection $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow G$ as before. Then, for every prime $p$, the local embedding problem (2.4) is solvable.
@@
First suppose that $p$ is unramified in $L$. In this case, $G_{p}$ is cyclic, generated by the Frobenius element. Let $\mathbb{Q}_{p}^{\operatorname{nr}}$ be the maximal unramified extension of $\mathbb{Q}_{p}$ and note that $\operatorname{Gal}(\mathbb{Q}_{p}^{\operatorname{nr}}/\mathbb{Q}_{p})$ is isomorphic to $\widehat{\mathbb{Z}}$. We seek a lift $\widetilde{\varphi}_{p}\mathrel{\mathop{\ordinarycolon}}\widehat{\mathbb{Z}}\rightarrow\widetilde{G}_{p}$ of $\varphi_{p}$. In order to show that $\widetilde{\varphi}_{p}$ exists, it suffices to prove that $\widetilde{G}_{p}$ is cyclic. Indeed, $G_{p}$ is cyclic, in particular, abelian. Hence $\mathbb{Z}/\ell\mathbb{Z}$ contains $\widetilde{G}_{p},\widetilde{G}_{p}]$. If $\mathbb{Z}/\ell\mathbb{Z}=[\widetilde{G}_{p},\widetilde{G}_{p}]$, then, $\mathbb{Z}/\ell\mathbb{Z}$ is contained in $\Phi(\widetilde{G}_{p})=[\widetilde{G}_{p},\widetilde{G}_{p}]\widetilde{G}_{p}^{\ell}$. Therefore, there is a surjection $\widetilde{G}_{p}/\left(\mathbb{Z}/\ell\mathbb{Z}\right)\twoheadrightarrow\widetilde{G}_{p}/\Phi(\widetilde{G}_{p})$. In particular, we find that $\widetilde{G}_{p}/\Phi(\widetilde{G}_{p})$ must be cyclic. It follows that $\widetilde{G}_{p}$ must be cyclic as well. This contradicts the assumption that $\widetilde{G}_{p},\widetilde{G}_{p}]=\mathbb{Z}/\ell\mathbb{Z}$. Thus, $\widetilde{G}_{p}$ is abelian. Since $G_{p}$ is cyclic and the sequence (2.4) is non-split, this forces $\widetilde{G}_{p}$ to be cyclic. Therefore, $\varphi_{p}\mathrel{\mathop{\ordinarycolon}}\widehat{\mathbb{Z}}\rightarrow G_{p}$ lifts to a map $\widetilde{\varphi}_{p}\mathrel{\mathop{\ordinarycolon}}\widehat{\mathbb{Z}}\rightarrow\widetilde{G}_{p}$.